# HCT-QA: A Benchmark for Question Answering on Human-Centric Tables

### v2

@@ C.1 Generating process
C.2 Semantic templates
Figure 30: Semantics for generating synthetic tables and questions. Examples of $T_{REL}$ (a) and a corresponding $T_{HCT}$ (b) related to the amount of food import-export in some place. (c) JSON format to describe Food Import-Export semantic terms used to generate these tables (see Section C.3) and the related questions (see Section C.4). Not all semantics terms are used in $T_{REL}$ and $T_{HCT}$.
In this section, all terms refer to Figure 30.
Our synthetic data relies on semantic terms naming the dataset’s attributes and values. A synthetic $T_{HCT}$ (Figure 30b) is obtained by pivoting a relational table $T_{REL}$ (Figure 30a) composed of at least two nominal attributes and a single numerical attribute always named Value.
There are two types of nominal attributes:
  • Hierarchical attributes like food Category (e.g.Meat, Cereals, Dairy, Beverage) and food Item (e.g.Cream, Soda, Milk) are dependent on each other. If both attributes appear in the HCT, the food Item must be displayed at a lower level than the food Category, i.e. closer to the inner part of the HCT, and all must appear together either in row headers or column headers.
Semantics are encoded in a JSON object (Figure 30c) with two main keys: data and values.
  • data refers to a list of attributes (e.g.ImportExport, Years, Quarters) or groups of attributes (e.g.Food), each described with three keys:
    • code is the primary key used in table and question templates to identify the independent attributes (e.g.ImportExport or Years) and the hierarchical groups of attributes (e.g.Food).
    • names is the list of attribute names displayed in $T_{REL}$ column headers but typically hidden (implicit) in $T_{HCT}$ and $Q_{NL}$. It is a list with a single name for independent attributes, or with as many names as levels for hierarchical groups of attributes (e.g.Food), ordered from top (outer) to bottom (inner) levels (e.g.Category, Item).
    • values is the list of values (e.g.2015, 2016... for the attribute Year). It can also be the hierarchy of values of a group of attributes. For instance, in Food group, Meat is a value of the attribute Category, and Beef, Chicken…, are values of the attribute Item, whose parents are Meat in the Food hierarchy. Attribute values are always displayed as row or column headers in $T_{HCT}$.
  • values refers to the numerical values to be generated in both the core part of $T_{HCT}$ and the Value column of $T_{REL}$. This object contains shortcut keys (e.g.realUnit or intSigned) to generate specific ranges and types of numbers when called in table templates. The value associated with such shortcuts must be a list of two numbers $[m,M]$ representing the minimum $m$ and maximum $M$ possible values. If any of these min and max is a decimal number, the generated value will be a real number with two decimal digits drawn uniformly in this interval. If both min and max are integers, the generated value will be an integer drawn uniformly in this interval. The generator will ensure no replica in the generated values, reals or integers, possibly overflowing the max value if asked to generate more numeric values than the range allows.
Customization of semantic data
All terms in semantic data are customizable except the JSON keys data, code, names, and values, following the rules above. The number of decimals of generated real numbers is specified by NUM_DECIMAL_DIGITS_REAL_FORMAT in config.R. These semantic data are stored in the file PARAM_semantics.json.
C.3 Table templates
Figure 31: Example of a synthetic table template JSON file described in Section C.3, used to generate sets of Food Import/Export HCT and relational tables. This template is stored in the file PARAM_tableTemplates.json
The generic table templates described in the file PARAM_tableTemplate.json are used to generate a (large) set of $T_{REL}$ and $T_{HCT}$ tables from the semantic data (See Section C.2). All terms below refer to Figure 31.
This template contains the following keys:
  • tables is a list of table content specifications (values, headers, and aggregates) with coherent semantics. Each initial table template in the list has the following keys:
    • values is either a shortcut name referring to one of the keys (e.g.realUnit) stored in values in the semantic data (see Figure 30), or directly an interval of the form $[m,M]$ following the description of the main key values in Section C.2
[…]
    • rowCodes is a list of one or more of the data code keys (e.g.Food or Quarters) listed in the semantic data. The generator will pick among these codes to generate the rows of $T_{HCT}$, and one part of the nominal attributes of $T_{REL}$.
    • rowSamples is a list containing as many intervals $[m,M]$ as there are codes listed in rowCodes, in the same order. Each interval specifies the minimum $m$ and maximum $M$ number $n$ of values to select among the list $L$ of values of the specified attribute. The generator first picks $n$ at random in $[m,M]$, then selects $n$ consecutive values in $L$ starting at a random position in $L$ but ensuring there are always $n$ values in the result if the list $L$ is long enough, or return $L$ itself otherwise. For instance, let’s consider an independent attribute Years with a list of values: $L=[2001,2002,2003,2004,2005,2006]$, if we set rowSamples to $[1,5]$, and the generator randomly picks $n=3$, it then picks a random index $i$ between $1$ and $4$, e.g.i=2i=2, to generate the final subset $[2002,2003,2004]$ of years to use in the table. For hierarchical attributes, the same process applies recursively to all levels of the hierarchy, starting from the top level. If the interval is set to $[0]$ instead, all the values of the corresponding independent attribute are generated, or all the values at all hierarchy levels are generated for hierarchical attributes. Note that the col_row_levels parameter limits the depth of row and column headers, so the attributes finally displayed in $T_{HCT}$.
    • colSamples is the same as rowSamples but for columns.
    • agg_name1 is the name of the aggregate to display in the table.
These generic table templates PARAM_tableTemplate.json are used with the semantic data PARAM_semantics.json to generate individual templates using the script 1_geneAllTablesJSONfromPatterns.R. The individual templates generated and described below are stored in the file PARAM_tablesToGenerate.json.
Customization of generic table templates
These generic table templates are customizable except for all the JSON keys and following the rules above.
An individual table template is a JSON object (Figure 32) instantiating one combination of the parameters shuffle, col_row_levels, col_row_name_pos, col_row_agg_pos, and row_format common to all initial table templates, and set in the generic table template. It contains the following keys:
  • replica is identical to replica set in the generic table template.
  • agg_fun1 is identical to agg_fun1 set in the generic table template.
  • values is identical to values set in the generic table template.
Two examples of individual templates with two of the $5$ sampled $T_{REL}$ and corresponding $T_{HCT}$ are shown in Figure 32.
Customization of individual table templates
There is no possible customization of these individual table templates automatically generated from the generic table templates.
C.4 Question templates
Figure 33: Synthetic question templates for the Food Import/Export (a) and Number of Students (b) table semantics.
Figure 33 shows the natural language question templates related to ”Food import-export in tons” (a) and ”Number of students” (b) semantics and tables, as a JSON object with the following keys:
  • tableName is identical to the valueName key in the generic table template (see Section C.3).
@@ C.4 Question templates
  • template_N a list of templates for questions of type N (See the list of $15$ types in Section C.1).
@@
    • We use \$attribute_name to specify variables that will be automatically replaced by the generator with values from that attribute present in the table instance (e.g.\$Category could be replaced by Meat, Dairy, Cereals, or Beverage). These attribute_name must appear in the rowColCodes list for independent attributes, and in the simplifyNested list for groups of hierarchical attributes.
@@
    • In template $12$, top/bottom_\$TOPK will be replaced by an integer and either top or bottom automatically selected, e.g.”What are the top 5 amounts of Import of…
@@
    • We use _ to link words like of, in to their attribute. If an attribute is not present in an instance of a question, all the words connected to it with _\_ will be absent too (e.g.of_\$Import_Export).
@@
All $14$ templates are repeated with variations for the $7$ semantics in the PARAM_NLquestionTemplates.json file, guiding the creation of the same templates for new semantics. Note that the question $Q_{NL}15$ is automatically generated from templates $13$ and $14$.
@@
These question templates are customizable except for all the JSON keys and following the rules above.
The Figure 34 shows the matching between the synthetic question types and the question properties described in Section B.
Figure 34: Correspondence between synthetic question templates (types 11 to $15$ in columns) and question properties (rows). Cells are green when the property is present in the template type. $n$ is a numeric value that can vary for a given type.
C.5 Synthetic generation code
The list of codes and the instructions for installing and running them are given in https://github.com/shahmeer99/HCT-QA-Benchmark/blob/main/synthetic_data_generator/README_SYNTHETIC_GENERATOR.md.
Appendix D Additional Experiments and Details
D.1 Model Details
Table 8 contains all the information about the models used in our experiments.
Table 8: Summary of Evaluated Language Models
Model Name Provider Size
TEXT-ONLY LLMs
Large Models – 100B+ (Closed)
gpt-o1-mini OpenAI 100B+
gpt-4o OpenAI 100B+
gpt-35 OpenAI 100B+
Medium Models – 27B-70B (Open)
Qwen2.5-72B-Instruct Alibaba 72B
Llama-3.1-70B-Instruct Meta 70B
Llama-3.3-70B-Instruct Meta 70B
gemma-2-27b-it Google 27B
gemma-3-27b-it Google 27B
Mixtral-8x7B-Instruct-v0.1 MistralAI 56B
Small Models – 3B-12B (Open)
Meta-Llama-3-8B-Instruct Meta 8B
Llama-3.1-8B-Instruct Meta 8B
Qwen1.5-7B-Chat Alibaba 7B
Qwen2-7B-Instruct Alibaba 7B
Qwen2.5-7B-Instruct Alibaba 7B
gemma-2-9b-it Google 9B
gemma-3-12b-it Google 12B
Mistral-7B-Instruct-v0.3 MistralAI 7B
Mathstral-7B-v0.1 MistralAI 7B
FanarC QCRI 9B
aya-23-8B CohereForAI 8B
Phi-3.5-mini-instruct Microsoft 3.8B
Model Name Provider Size
VISION-TEXT VLMs/MLMs
Large Models – 100B+ (Closed)
gpt-4o-vision OpenAI 100B+
Small Models – 3B-12B (Open)
Pixtral-12B-2409 MistralAI 12B
InternVL2-4B SenseTime 4B
InternVL2_5-8B-MPO SenseTime 8B
Phi-3-vision-128k-instruct Microsoft 4.2B
Phi-3.5-vision-instruct Microsoft 4.2B
Idefics3-8B-Llama3 HuggingFace/Inference.ai 8B
Qwen2-VL-7B-Instruct Alibaba 7B
llava-1.5-7b-hf LMSYS Org (LLaVA) 7B
llava-v1.6-mistral-7b-hf MistralAI (LLaVA) 7B
llava-v1.6-vicuna-7b-hf LMSYS Org (LLaVA) 7B
Molmo-7B-D-0924 ImitateML 7B
paligemma2-10b-ft-docci-448 Google 10B
D.2 Additional Trend Analysis & Observations
(RQ6) Does model verbosity affect the output correctness?
@@ D.2 Additional Trend Analysis & Observations
Our initial hypothesis was that if models simply take the kitchen-sink approach and output more values in their answers then they would have a higher probability of matching values in the ground truth and thereby inflate their recall scores. As shown in Figure 35, there is no clear correlation between answer length and recall. Most models, including the best performing ones, seem to be clustered around the middle on the verbosity scale. The average number of values in the ground-truth answers for all queries is  4, yet most models tend to provide between 2-3 values in their answers. This suggests that models may in fact be conservative in their answer verbosity as which is antithetical to the overly verbose nature of these generative models.
@@
Figure 35: Recall Score vs. Verbosity for Text-Only Models on Real World HCTs
(RQ7) How do question properties affect model performance?
The synthetic queries we generated have 15 types including various types of selection, projection, expression, and aggregation. The model performance per template type is shown in Figure 36. As expected, models tend to do best on selection queries be it single-cell or multi-cell selection. This is clear as templates 1, 2, and 3 have the highest average recall across all models. Models tend to do better on queries involving expressions on columns (templates 5, 14) than queries involving aggregations (template 7). In fact, models struggle the most with template 7 which requires doing an actual aggregation (rather than retrieving an existing one from the HCT). Models like LLaMA-3.3-70B significantly outperform even models like GPT-4o and GPT-35 on this template. It is also evident that models clearly have their strengths and weaknesses in terms of templates. GPT-4o gets 58% on template 11, whereas other models like LLaMA and Mistral do not even cross 10%. However, this dominance completely disappears in Templates 4, 7, and 9 where the GPT-4o fails to score more than 8% (compared to other models that get up to 42%). There are also unexpected trends across model generations. Gemma-2-27B scores a 52% recall on template 14 which is only topped by GPT-4o and Qwen2.5 72B, both of which are relatively larger models. However, on the exact same template (14), Gemma-3-27B scores exactly half of its predecessor with a recall score of 26%. This shows that newer generations of models are not necessarily improving their capabilities on this task across the board.
@@
Figure 36: Recall Score for Text-Only Models on all Synthetic HCTs per Question Template
(RQ8) How are model size and performance correlated?
It is clear that large closed-weight models like GPT are far superior to their smaller open-weight competitors. However, a fairer comparison to answer this question would be between the same model version but in different sizes. Figure 37 shows exactly that and we can see that larger models tend to do considerably better than their smaller versions. LLAMA-3.1 improved its recall score by nearly 25% with a 9x increase in parameter size. Similarly, Gemma-2 improves its recall score by 11% with a 3x parameter size increase. The improvement between Mistral-7B and Mixtral-8x7B is only 8% but this is most likely due to the fact that the latter is a Mixture-of-Experts model where not all experts (active parameters) are being used for this task. Overall, increase in parameter size does improve performance to a good degree. On average when we consider LLaMA3.1, Gemma-2, Gemma-3, and Qwen2.5 we see an approximate 4% increase in recall score per every 10B additional parameters.
Figure 37: Recall Score Across Size (Number of Parameters) within Model Families
(RQ9) In what cases do VLMs perform better than LLMs?
@@
Recent advances in the image understanding and reasoning capabilities of VLMs make them a suitable candidate for the task at hand. VLMs also allow to skip the extraction and conversion to CSV process as they can directly consume the images of the tables from their source documents. On average, VLMs score 25.5% recall whereas text-only LLMs score 25% as shown in (Table 1). GPT4o when prompted using the text-modality achieves an average recall of 61.6%, compared to 62.3% when using the image-modality. This illustrates the progress vision models have made in terms of document understanding and reasoning tasks such as this one. We have established that HCTs inherently have visual characteristics that are crucial to understand the content in the table. A lot of those visual clues are lost when the HCT is converted to text format (regardless of which text format its represened in). VLMs are able to overcome this information loss by taking advantage of those visual cues. This is further shown by the top performing small sized VLM model, Pixtral-12B with a recall score of 45.9% which is nearly 8% better than the best performing text-only LLM of that size range (Qwen2.5-7B with recall 38%). Furthermore, Microsoft’s Phi-3-Vision-128k (image-modality), which is only a 4.2B parameter model outperforms every text-only LLM in the small size category (many of which are 2x in size). These results clearly highlight the potential of VLMs in visual-reasoning tasks such as HCT-QA. Another interesting observation is that VLMs struggle considerably more than LLMs with larger HCTs. GPT4o performs almost identically in both the text and vision modalities except in the US CENSUS dataset which on average has 437 cells per table (nearly 3x more compared to the other datasets, see Table 5). This trend can also be seen in the smaller VLMs as most top performing models drop in performance when dealing with US CENSUS HCTs.
@@
Figure 38: LLMs vs. VLMs on all Real World HCTs - Recall Score
Regression analysis of table properties and mean recall score
(across all models) is shown in Table 9.
Table 9: Effect of table properties on Recall. Only properties with significant effect ($p<0.05$) are shown
Table Property Coef. (p-value)
Symmetric Row Nesting -0.103 (*0.016)
Asymmetric Row Nesting -0.142 (*0.001)
Symmetric Col. Nesting -0.095 (0.148)
Asymmetric Col. Nesting -0.127 (0.061)
Balanced Row Nesting -0.251 (0.073)
Balanced Col. Nesting -0.270 (*0.022)
Unbalanced Col. Nesting -0.315 (*0.009)
Global Column Aggreg. -0.035 (*0.028)
Implicit Col. Aggreg. 0.081 (*0.001)
Implicit Row Aggreg. -0.065 (*0.007)
Row Group Label -0.069 (*0.000)
D.3 Prompting Details & Experiments
@@
D.3.1 Prompts
@@
User Prompt for Inference Answer the question given the input table. Do not provide any explanations, equations, code, or text explaining intermediate steps in figuring out the answer. Return only the final answer itself. # INPUTS: Table: ‘{table}‘ Question: ‘{quesiton}‘ # OUTPUT: Answer:
D.3.2 0-shot vs. 1-shot prompting
The majority of the models we used in our experiments can handle context lengths of 8,000+ tokens. This means that it is possible to do 1-shot prompting with a full table-question-answer triplet as the shot. This approach could help leverage in-context learning to improve the model performance. We experimented on a small set of 1,000 queries using Real World HCTs from all four data sources. Our 1-shot prompt consisted of a fixed example for all the questions in the test set. We avoided testing the models given an example question on the same HCT the test question was on as that reflects a very unrealistic scenario. The results are shown in Table 10. We can see that adding the 1-shot example does not significantly affect the recall score as for gemma-2-9b-it it only increases by 1.3% and for Llama-3.1-8B-Instruct it actually decreases by less than 1%. Upon manual inspection, we found that the models were able to follow the output format stated in the prompt (values split by ‘ || ‘ and no explanations) in the 0-shot scenario almost perfectly. As the 1-shot example was on a different table, it provided no ”learning” for the model to answer the test question it was being asked. Due to these results, our remaining experiments use 0-shot prompting.
Table 10: 0-shot vs 1-shot Prompting - Recall Score
@@
D.4 Evaluation Metrics
We experimented using Meta-Llama-3-70B as judge model to gauge the correctness of a model’s answer. The prompt given to the judge model outlined a clear set of instructions: (1) minor spelling mistakes or typos in textual values can be ignored, (2) rounding numerical values is fine if done reasonably, (3) question must be answered fully in order to be ”correct”. We compared this model’s ”correct” labels to our CC Score (Complete Containment, which we defined as 1 when Recall Score == 1, else 0). The comparison is shown in Figure 39. It is evident that the model’s judgment and the CC score closely align with each other. Since running an LLM-as-a-Judge is expensive in terms of computational resources, we opt to use Recall, Precision, and CC Score as our evaluation metrics.
Figure 39: CC Score vs. LLM-as-a-Judge Score on Real World HCTs
D.5 VLMs vs LLMs
We can see from the results in Table 1 that all VLMs are sensitive to table size. HCTs arXiv and Pak have relatively lower number of cells and number of words on average and vision models tend to do the best on these tables. It is evident that all vision models struggle with US Census tables (which are the largest tables on average with good margin). Gpt-4o scores 39.6% recall on US Census which is less than majority of medium sized text-only models which are considerably smaller in size.

### v3

@@ B-D Natural Language question templates
Figure 36: Synthetic natural question templates for the Accidents (a), Constructions (b), Food (c), and Graduations (d) domains.
Figure 37: Synthetic natural question templates for the Pollution (a), Students (b), and Weather (c) domains.
Figures 36 and 37 show the natural language question templates related to the seven generated HCT domains, as JSON objects with the following keys:
  • tableName is identical to the valueName key in the generic table template (see Section B-C).
@@
  • template_N a list of templates for questions of type N (See the list of $15$ $Q_{NL}$ templates above).
@@
    • We use \$attribute_name to specify variables that will be automatically replaced by the generator with values from that attribute present in the table instance (e.g., \$Category could be replaced by Meat, Dairy, Cereals, or Beverage). These attribute_name must appear in the rowColCodes list for independent attributes, and in the simplifyNested list for groups of hierarchical attributes.
@@
    • In template $12$, top/bottom_\$TOPK will be replaced by an integer and either top or bottom automatically selected, e.g., ”What are the top 5 amounts of Import of…
@@
    • We use _ to link words like of, in to their attribute. If an attribute is not present in an instance of a question, all the words connected to it with _ will be absent too (e.g., of_\$Import_Export).
@@
All $14$ templates are repeated with variations for the $7$ domains in the PARAM_NLquestionTemplates.json file, guiding the creation of the same templates for new domains. Note that the question $Q_{NL}15$ is automatically generated from templates $13$ and $14$.
@@
Customization of question templates These question templates are customizable except for all the JSON keys and following the rules above.
Appendix C Additional Experiments and Details
C-A Additional Trend Analysis & Observations
@@
(RQ8) How does model verbosity affect the output correctness? Our initial hypothesis was that if models simply take the kitchen-sink approach and output more values in their answers then they would have a higher probability of matching values in the ground truth and thereby inflate their recall scores. As shown in Figure 38, there is no clear correlation between answer length and recall. Most models, including the best performing ones, seem to be clustered around the middle on the verbosity scale. The average number of values in the ground-truth answers for all queries is  4, yet most models tend to provide between 2-3 values in their answers. This suggests that models may in fact be conservative in their answer verbosity as which is antithetical to the overly verbose nature of these generative models.
@@
Figure 38: Recall Score vs. Verbosity for Text-Only Models on Real World HCTs.
(RQ9) How do model perform on different synthetic QA templates? The synthetic queries we generated have 15 types including various types of selection, projection, expression, and aggregation. The model performance per template type is shown in Figure 39.
@@
Figure 39: Recall Score for Text-Only Models on all Synthetic HCTs per Question Template.
As expected, models tend to do best on selection queries be it single-cell or multi-cell selection. This is clear as templates 1, 2, and 3 have the highest average recall across all models. Models tend to do better on queries involving expressions on columns (templates 5, 14) than queries involving aggregations (template 7). In fact, models struggle the most with template 7 which requires doing an actual aggregation (rather than retrieving an existing one from the HCT). Models like LLaMA-3.3-70B significantly outperform even models like GPT-4o and GPT-35 on this template. It is also evident that models clearly have their strengths and weaknesses in terms of templates. GPT-4o gets 58% on template 11, whereas other models like LLaMA and Mistral do not even cross 10%. However, this dominance completely disappears in Templates 4, 7, and 9 where the GPT-4o fails to score more than 8% (compared to other models that get up to 42%). There are also unexpected trends across model generations. Gemma-2-27B scores a 52% recall on template 14 which is only topped by GPT-4o and Qwen2.5 72B, both of which are relatively larger models. However, on the exact same template (14), Gemma-3-27B scores exactly half of its predecessor with a recall score of 26%. This shows that newer generations of models are not necessarily improving their capabilities on this task across the board.
@@
(RQ10) In what cases do VLMs perform better than LLMs? Recent advances in the image understanding and reasoning capabilities of VLMs make them a suitable candidate for the task at hand. VLMs also allow to skip the extraction and conversion to CSV process as they can directly consume the images of the tables from their source documents. On average, VLMs score 25.5% recall whereas text-only LLMs score 25% as shown in (Table IV). GPT4o when prompted using the text-modality achieves an average recall of 61.6%, compared to 62.3% when using the image-modality. This illustrates the progress vision models have made in terms of document understanding and reasoning tasks such as this one. We have established that HCTs inherently have visual characteristics that are crucial to understand the content in the table. A lot of those visual clues are lost when the HCT is converted to text format (regardless of which text format its represented in). VLMs are able to overcome this information loss by taking advantage of those visual cues. This is further shown by the top performing small sized VLM model, Pixtral-12B with a recall score of 45.9% which is nearly 8% better than the best performing text-only LLM of that size range (Qwen2.5-7B with recall 38%). Furthermore, Microsoft’s Phi-3-Vision-128k (image-modality), which is only a 4.2B parameter model outperforms every text-only LLM in the small size category (many of which are 2x in size). These results clearly highlight the potential of VLMs in visual-reasoning tasks such as HCT-QA. Another interesting observation is that VLMs struggle considerably more than LLMs with larger HCTs. GPT4o performs almost identically in both the text and vision modalities except in the US CENSUS dataset which on average has 437 cells per table (nearly 3x more compared to the other datasets. This trend can also be seen in the smaller VLMs as most top performing models drop in performance when dealing with US CENSUS HCTs.
@@
Figure 40: LLMs vs. VLMs on all Real World HCTs — Recall Score.
C-B Prompting Details & Experiments
@@
C-B1 Prompts
Prompts are used with LLMs to give them context and role, and to format the results to allow automatic evaluation.
@@ C-B1 Prompts
User Prompt for Inference Answer the question given the input table. Do not provide any explanations, equations, code, or text explaining intermediate steps in figuring out the answer. Return only the final answer itself. # INPUTS: Table: ‘{table}‘ Question: ‘{question}‘ # OUTPUT: Answer:
C-B2 0-shot vs. 1-shot prompting
The majority of the models we used in our experiments can handle context lengths of 8,000+ tokens. This means that it is possible to do 1-shot prompting with a full table-question-answer triplet as the shot. This approach could help leverage in-context learning to improve the model performance. We experimented on a small set of 1,000 queries using Real World HCTs from all four data sources. Our 1-shot prompt consisted of a fixed example for all the questions in the test set. We avoided testing the models given an example question on the same HCT the test question was on as that reflects a very unrealistic scenario. The results are shown in Table VIII. We can see that adding the 1-shot example does not significantly affect the recall score as for gemma-2-9b-it it only increases by 1.3% and for Llama-3.1-8B-Instruct it actually decreases by less than 1%. Upon manual inspection, we found that the models were able to follow the output format stated in the prompt (values split by ‘ || ‘ and no explanations) in the 0-shot scenario almost perfectly. As the 1-shot example was on a different table, it provided no ”learning” for the model to answer the test question it was being asked. Due to these results, our remaining experiments use 0-shot prompting.
TABLE VIII: 0-shot vs 1-shot Prompting - Recall Score
@@
C-C Evaluation Metrics
We experimented using Meta-Llama-3-70B as judge model to gauge the correctness of a model’s answer. The prompt given to the judge model outlined a clear set of instructions: (1) minor spelling mistakes or typos in textual values can be ignored, (2) rounding numerical values is fine if done reasonably, (3) question must be answered fully in order to be ”correct”. We compared this model’s ”correct” labels to our CC Score (Complete Containment, which we defined as 1 when Recall Score == 1, else 0) on a subset of 500 real-world HCT QA pairs. The comparison is shown in Figure 41. It is evident that the model’s judgment and the CC score closely align with each other. Since running an LLM-as-a-Judge is expensive in terms of computational resources, we opt to use Recall, Precision, and CC Score as our evaluation metrics.
Figure 41: CC Score vs. LLM-as-a-Judge Score on Real World HCTs.
C-D VLMs vs LLMs
We can see from the results in Table IV that all VLMs are sensitive to table size. HCTs arXiv and Pak have relatively lower number of cells and number of words on average and vision models tend to do the best on these tables. It is evident that all vision models struggle with US Census tables (which are the largest tables on average with good margin). Gpt-4o scores 39.6% recall on US Census which is less than majority of medium sized text-only models which are considerably smaller in size.